Ask your own question, for FREE!
Mathematics 7 Online
OpenStudy (daniellelovee):

A pond is freshly stocked with 15 brown trout and 9 lake trout. The first fisherman in the area catches and releases a trout. He catches another trout a while later. What is the probability that the fisherman caught brown trout each time?

OpenStudy (daniellelovee):

@Directrix

OpenStudy (daniellelovee):

@dan815

Directrix (directrix):

15 brown trout 9 lake trout 15 + 9 = 24 total trout 15/ 24 * 15/24 = 5/8 * 5/8 = 25 / 64 Probability = 25/64 The fish is released so the probability is the same both times. @Daniellelovee Check the calculations for accuracy.

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
17 hours ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
20 hours ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Arizona: does anyone know the percentage rate of 30,000 and 40,000
1 week ago 0 Replies 0 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!